OpenStudy (anonymous):

Solve for V 4 = − 2 / 9 v Simplify your answer as much as possible.

OpenStudy (aaronandyson):

wait

OpenStudy (aaronandyson):

It's 9v or just v?

OpenStudy (anonymous):

just v

OpenStudy (aaronandyson):

\[4 = \frac{ -2 }{ 9 } \times v?\]

OpenStudy (aaronandyson):

@NadiaElhaj_1

OpenStudy (anonymous):

it just shows it like 4 = - 2/9 v it doenst show the *v and the negative isn't just on the 2 its in the middle of 3/9

OpenStudy (aaronandyson):

I'll assume it as 9v then?

OpenStudy (aaronandyson):

so 4 times 9v = -2 36v = -2 v = 36/-2 v = -18

OpenStudy (misty1212):

\[\huge 4=-\frac{2}{9}v\] like that ?

OpenStudy (misty1212):

or \[\huge 4=-\frac{2}{9v}\] there is a big difference

OpenStudy (anonymous):

the first one @misty1212

OpenStudy (anonymous):

@AaronAndyson can u help please

syed98 (syedmohammed98):